A rocket is fired at a ground level and climbs at a constant angle of 75degrees. what is the altitude of the rocket when measured from a distance of 10,000 feet from the launch spot?
I know to start this problem you need a right triangle and the base of the triangle is 10,000 feet but im not sure what else please help very stuck

Answer by fcabanski(874)

Draw the triangle. The 75 degree angle is between the base and the hypotenuse. The h is the side opposite that angle, while the 10,000 base is the side adjacent to that 75 degree angle.


This is a job for tan = opp/adj = h/10000.


tan75 = 3.7321 = h/10000 (multiply both sides by 10000) h = 37321 feet.


I can help you understand this with one on one online tutoring. It's one thing to follow an example, but even better if you understand the concept. My email address is the web site address shown in my profile.